### v1

@@ Abstract
The number of objects in orbit is rapidly increasing, primarily driven by the launch of megaconstellations, an approach to satellite constellation design that involves large numbers of satellites paired with their rapid launch and disposal. While satellites provide many benefits to society, their use comes with challenges, including the growth of space debris, collisions, ground casualty risks, optical and radio-spectrum pollution, and the alteration of Earth’s upper atmosphere through rocket emissions and reentry ablation. There is substantial potential for current or planned actions in orbit to cause serious degradation of the orbital environment or lead to catastrophic outcomes, highlighting the urgent need to find better ways to quantify stress on the orbital environment. Here we propose a new metric, the CRASH Clock, that measures such stress in terms of the time it takes for a catastrophic collision to occur if there are no collision avoidance manoeuvres or there is a severe loss in situational awareness. Our calculations show the CRASH Clock is currently 2.8 days, which suggests there is now little time to recover from a wide-spread disruptive event, such as a solar storm. This is in stark contrast to the pre-megaconstellation era: in 2018, the CRASH Clock was 121 days.
@@ 1 Introduction
The use of KEIs can further help to provide a suite of metrics for exploring when nonlinearities could occur or where there could be system-wide effects (such as across LEO). As an example, the concept of Kessler-Cour-Pallais Syndrome (KCPS) [Kessler1978] is made evident by KEIs (e.g., critical densities or collision area). KCPS is a condition in which the growth of collisional debris (and surface area) on orbit outpaces the debris’ removal through atmospheric drag, causing a collisional runaway. However, despite the use of the term “runaway”, the initial phase of KCPS (which some argue we have already entered, see e.g. [Kelvey2024]) is characterized by slow growth of debris, taking decades to centuries to develop. Such long timescales create a challenge for using the idea of KCPS alone as a policy tool.
@@
With this in mind, we introduce the Collision Realization And Significant Harm (CRASH) Clock, a KEI that evaluates the stress on the orbital environment. Many simulations predicting collision rates on orbit assume perfect collision avoidance for all manoeuvrable payloads during their operational lifetimes (e.g. Section 7.2 of [ESA2025], [Letizia2023]). In contrast, the CRASH Clock uses the known distribution of resident space objects (RSOs: active and derelict satellites, debris, and rocket bodies) to determine how quickly we could expect a collision if collision avoidance manoeuvres were to suddenly stop or if there was a severe loss in situational awareness (such as due to a major solar storm or catastrophic software issue). It is a measure, in part, of the degree to which the orbital environment is a house of cards.
@@ 2 Results
Using this density distribution, we can calculate expectations for close encounter rates among RSOs, as well as collisions under the assumption of no collision avoidance manoeuvres. The satellite-satellite encounter/collision rate is approximated by
$\Gamma_{\rm sat}=\int_{V}n_{\rm sat}^{2}A_{\rm col}\bar{v}_{r}\,dV,$ (1)
@@
where $n_{\rm sat}$ is the number density of satellites at some altitude, $A_{\rm col}$ is the collision cross-section of the satellites, $\bar{v}_{r}$ is the typical relative collision speed, and $dV$ is the spherical volume element. This collision rate can naturally be extended to include combinations of various RSOs (e.g., debris-satellite, debris-debris), as well as non-trivial distributions of collisional area. We also assume that the satellites are randomly distributed within a shell.
@@
To demonstrate the consequences of Figure 1, we can integrate Equation 1 over a limited volume. As an illustrative example, let us integrate over the peak in the density distribution centred on 550 km altitude, assuming a uniform spherical shell. We approximate the shell as being 30 km thick from the peak’s width, and assume an average collision cross-section of $A_{\rm col}\approx 300~\rm m^{2}$, which we consider to be conservative. As discussed in Methods, we use $\bar{v}_{r}\approx 10~\rm km~s^{-1}$. The density appears to peak at $n_{\rm sat}\approx 3\times 10^{-7}~\rm km^{-3}$, but upon closer inspection we find that the density finely oscillates around $n_{\rm sat}\approx 2\times 10^{-7}~\rm km^{-3}$, so we use this latter value as the average density in the shell.
@@
For the given altitude, the single shell collision rate is thus $\Gamma_{\rm ss}\approx 2.2\times 10^{-6}~\rm s^{-1}$ or about $0.19\rm~d^{-1}$. Defining the single shell expectation time between collisions to be $\tau_{\rm ss}=1/\Gamma_{\rm ss}$, we find $\tau_{\rm ss}\approx 5.3~\rm d$. This means that the timescale for a 50% chance of one or more collisions is about $t\approx 3.7\rm~d$ in that shell alone, assuming a Poisson process, i.e. probability $P=1-\exp(-t/\tau_{\rm ss})$. This result, while approximate, emphasizes that the collision probability for satellites on orbit is substantial without frequent collision avoidance.
@@
Although spherical shells around Earth represent extremely large volumes and the instantaneous volume occupied by satellites is small, LEO satellites orbit the Earth in approximately 90 minutes depending on the altitude. As a result, these satellites quickly explore their mutual interaction possibilities. Collision avoidance manoeuvres are essential in dense satellite shells and successful repeated execution of these manoeuvres is the only reason why there has not been a recent major satellite-satellite collision as orbital densities continue to increase.
@@
With this in mind, the corresponding encounter time within each spherical shell is $\tau_{1,h}=1/\Gamma_{1,h}$. The blue curve shows this result in Figure 2, using the density distribution from Figure 1. In the densest part of Starlink’s 550 km orbital shell, we expect close approaches ($1~\rm{km}$) every 11 minutes in that shell alone.
@@
The time between <1<1 km close encounters for all of LEO is found by summing over the rates, with $\tau_{1}=\big(\sum_{h}\Gamma_{1,h}\big)^{-1}$. For all objects, Equation 1 gives $\tau_{1}=20$ seconds. This is dominated by encounters involving at least one satellite (i.e. Sat-RSO encounters), with Starlink satellites making up the majority of the satellite population. For Sat-RSO encounters, $\tau_{1}\approx 22~\rm{s}$, while for Starlink-RSO, $\tau_{1}\approx 27~\rm{s}$.
According to the most recent SpaceX biannual report, Starlink satellites made 144,404 collision avoidance manoeuvres in the period between 1 Dec 2024 and 31 May 2025 [SpaceX2025], averaging to 41 manoeuvres per satellite per year, or one collision avoidance manoeuvre every 1.8 minutes across the whole megaconstellation. Keeping in mind that our encounter distance does not directly translate to a collision probability threshold for manoeuvres, the Starlink manoeuvre time is reasonably similar to our simulation results (see Methods and following discussion) and analytic calculation (1.1 min and 27 s, respectively).
@@ 2.1 The CRASH Clock
Taking this one step further, we can approximate the time for an actual collision on orbit, which we propose is an excellent measure of the stress in orbital space: the CRASH Clock. For this calculation, we must make some assumptions about typical collision cross sections. We use our previously calculated RSO number density distributions as described in Methods, but keep the densities separated into categories (i.e. the number density in a given shell is the sum of $n_{\rm debris}$, $n_{\rm sat}$, etc.). Each interaction combination is then assigned a collision cross section (described in Methods). The total collision rate for a given altitude is then
$\Gamma_{h}=\bar{v}_{r}V_{h}\sum_{ij}n_{i}n_{j}A_{ij}^{\rm col}$ (2)
@@
for object types $i$ and $j$, shell volume $V_{h}$ and average relative speed $\bar{v}_{r}=\bar{v}_{r}(h)$. The physical collision time is found by summing the rates over all of LEO. For our June 2025 catalogue, this is about $\tau_{\rm col}=2.8\rm~d$ – the CRASH Clock value for that date. For comparison, the no-manoeuvre collision time for Starlink-RSO collisions is about $3.3\rm~d$. Said differently, if collision avoidance manoeuvres were to suddenly stop, and with our assumptions of collision cross-sections, within 24 hours there is a 30% chance of a collision between two catalogued RSOs and a 26% chance of a collision involving a Starlink satellite, assuming a Poisson process. Such collisions would be catastrophic, causing a major debris-generating event with high likelihood of secondary and tertiary collisions due to high orbital densities and local collision areas. We repeat this calculation using the LEO TLE catalogue from 1 January 2018, providing a reference value from prior to the megaconstellation era. In stark contrast, the CRASH Clock was 121 days, corresponding to a less than 1% chance of a collision happening between two RSOs within 24 hours under no-manoeuvre conditions.
@@
The fidelity of our analytic calculations is compared to direct N-body simulations (see Methods). Simulations were initialized using TLEs for two different epochs: 1 January 2018, and 25 June 2025. The red curves in Figure 2 show the average time between conjunctions less than $1\,{\rm km}$ as a function of altitude, which shows reasonable agreement with the analytical estimate. Recall that the analytic simulations assume random orbital configurations within each shell, while the simulations use the actual orbits. At least for counting close conjunctions in a statistical sense, our assumption of a random distribution does not appear to be a major limitation of the analytic approach. The left-hand panel of Figure 2 shows the 2018 distribution, the right-hand shows 2025. We see that the minimum time between conjunctions has decreased by about two orders of magnitude between 2018 and 2025.
@@
Both the simulations and the analytic calculations are intended to represent collision statistics for a snapshot in time, akin to a steady state system. That is, we make our calculations using only a chosen TLE catalogue epoch and do not incorporate evolution of the objects on orbit through launches, active debris removal, or PMD. We note that our simulations also do not incorporate atmospheric drag, which would serve to clear out debris at lower altitudes, decrease satellite altitudes, and potentially increase the rate at which satellite orbits would randomize without station-keeping efforts in place.
@@
Additional simulation runs were carried out on these populations using a much smaller time-step (see Methods) and a conjunction distance threshold of $100\,{\rm m}$ to search for conjunction events that could give rise to a collision. For the 25 June 2025 TLEs, we see a $100\,{\rm m}$ conjunction frequency of approximately 1 encounter every 56 minutes, within a factor of two of the analytic estimate (expecting approximately 1 encounter every 33 minutes for distances less than $100\,{\rm m}$). For this particular simulation initialization, a conjunction less than 30 metres occurs in the first 3 hours between debris and a satellite. We note that there is some difficult-to-characterize uncertainty in these numbers because we are using only a single simulation initialization. We have summarized all of our results for both TLE catalogue epochs in Table 1.
@@ 3 Discussion
In an effort to better characterize this stress, for both short and long-term consequences, we recommend the CRASH Clock, which is the expectation time for a collision if all satellite manoeuvres were to suddenly stop. It is a KEI calculated using the methods above, and provides an immediate assessment, as well as historical record, of the stress on the orbital environment. It further provides a measure of the typical time between system-wide loss of control in LEO and a probable catastrophic collision.
We find that our simulations and analytic calculation yield consistent results to within a factor of two. For the purpose of determining the CRASH Clock value, we recommend using the analytic method. There are several reasons for this. First, the analytic approach is an accessible calculation that only requires a TLE catalogue. Second, it avoids differences in the results that could arise from the choice of numerical methods. Third, simulation techniques could suffer from substantial statistical variation, potentially requiring many simulation initializations. Indeed, such variation can already be seen in the simulation results presented here.
Currently, the CRASH Clock in LEO is 2.8 days. Yet, the probabilistic nature of collisions means they could happen much sooner (again in the absence of manoeuvres). Recall that our example numerical simulations showed a debris-satellite conjunction of less than 30 metres in the first 3 hours of simulation time.
@@
A CRASH Clock value of less than three days is already a reason for concern, as major solar storms, such as the May 2024 Gannon storm, can have lingering impacts for the satellite population. For example, in the three days during and after this storm, more than half of all satellites (mostly Starlinks) manoeuvred due to increased atmospheric drag and to avoid subsequent collisions as a result of all the manoeuvres [Parker2024]. In such conditions, positional uncertainties can easily become as high as several km [Parker2024_2], making collision avoidance manoeuvres extremely uncertain. Moreover, while the May 2024 storm was the strongest geomagnetic storm in decades, The Great Geomagnetic Storm of 1859 was at least twice as intense [DSTDATA, CLIVER2013]. That event was characterized by two strong storms within a few days of each other, with the latter commonly known as the Carrington Event (September 1859). The storm peaks lasted for several hours, while the storm durations were for a day and several days, respectively [GREEN2006].
The number of collision avoidance manoeuvres made by Starlink has historically been doubling every six months [Pultarova]. Each manoeuvre creates uncertainty in the estimated satellite positions for multiple days, with one study even finding inaccuracies immediately after the manoeuvre of up to 40 km [Pultarova2]. As the number of required manoeuvres continues to increase, temporary lapses in collision avoidance capabilities, whether that be from inaccurate orbital determination or even a small miscommunication between operators in manoeuvre decision-making, will become increasingly catastrophic in their potential consequences. Indeed, in 2019 an ESA satellite was forced to manoeuvre out of the way of a Starlink satellite when a bug in SpaceX’s alert system prevented them from seeing an increased collision probability [boyle]. Before modern space traffic management policies, insufficient manoeuvre plan information between operators was the main cause that led to the Iridium-Cosmos collision of 2009 [shepperd].
@@
All the above discussion contextualizes how the the CRASH Clock can be used. It is not a strict limit - rather, it measures risk along a spectrum, with short CRASH Clock times representing a dangerous condition, moderate times representing a caution region, and long times indicating a healthy operational orbital environment. There is some degree of subjectivity in the boundaries of these “danger”, “caution”, and “safe” regions, but this makes the CRASH Clock adaptable without changing the underlying definition. As an illustrative example, and recalling our calculations above, our current CRASH Clock value of $\tau_{\rm col}=2.8~\rm d$ corresponds to a 30% probability of one or more collisions during a 24 hour period of no manoeuvre conditions. We consider this to be well within the “caution” region. We could further (and arbitrarily) define the “danger” region to be where the CRASH Clock value implies a 50% chance of at least one collision within 24 hours. To stay in the “caution” region and below this “danger” threshold, the CRASH Clock value would therefore need to be longer than 1.4 d.
In addition to the dangerously high collision risks calculated here, we are already experiencing disruption of astronomy [Lawrence2022], pollution in the upper atmosphere from increasingly frequent satellite ablation [Murphy2023], and increased ground casualty risks [Wright2025]. By these safety and pollution metrics, it is clear we have already placed substantial stress on LEO, and changes to our approach are required immediately.

### v2

@@ Abstract
The number of objects in orbit is rapidly increasing, primarily driven by the launch of megaconstellations, an approach to satellite constellation design that involves large numbers of satellites paired with their rapid launch and disposal. While satellites provide many benefits to society, their use comes with challenges, including the growth of space debris, collisions, ground casualty risks, optical and radio-spectrum pollution, and the alteration of Earth’s upper atmosphere through rocket emissions and reentry ablation. There is potential for current or planned actions in orbit to cause serious degradation of the orbital environment or lead to catastrophic outcomes, highlighting the urgent need to find better ways to quantify stress on the orbital environment. Here we propose a new metric, the CRASH Clock, that measures such stress in terms of the timescale for a possible catastrophic collision to occur if there are no satellite manoeuvres or there is a severe loss in situational awareness. Our calculations show the CRASH Clock is currently 5.5 days, which suggests there is limited time to recover from a wide-spread disruptive event, such as a solar storm. This is in stark contrast to the pre-megaconstellation era: in 2018, the CRASH Clock was 164 days.
@@ 1 Introduction
The use of KEIs can further help to provide a suite of metrics for exploring when nonlinearities could occur or where there could be system-wide effects (such as across LEO). As an example, the concept of Kessler-Cour-Pallais Syndrome (KCPS) [Kessler1978] is made evident by KEIs (e.g., critical densities or collision area). KCPS is a condition in which the growth of collisional debris (and surface area) on orbit outpaces debris removal through atmospheric drag, causing a collisional runaway. However, despite the use of the term “runaway”, the initial phase of KCPS (which some argue we have already entered, see e.g. [Kelvey2024]) is characterized by slow growth of debris, taking decades to centuries to develop. Such long timescales create a challenge for using the idea of KCPS alone as a policy tool.
@@
With this in mind, we introduce the Collision Realization And Significant Harm (CRASH) Clock, a KEI that evaluates the stress on the orbital environment. Many simulations predicting collision rates on orbit assume perfect collision avoidance for all manoeuvrable payloads during their operational lifetimes (e.g. Section 7.2 of [ESA2025], [Letizia2023]). In contrast, the CRASH Clock uses the known distribution of resident space objects (RSOs: active and derelict satellites, debris, and rocket bodies) to determine how quickly we could expect a collision if all manoeuvres were to suddenly stop or if there was a severe loss in situational awareness (such as due to a major solar storm or catastrophic software issue). It is a measure, in part, of the degree to which the orbital environment is a house of cards.
@@ 2 Results
Using this density distribution, we can calculate expectations for close encounter rates among RSOs, as well as collisions under the assumption of no manoeuvres. The satellite-satellite encounter/collision rate is approximated by
$\Gamma_{\rm sat}=\frac{1}{2}\int_{V}n_{\rm sat}^{2}A_{\rm col}\bar{v}_{r}\,dV,$ (1)
@@
where $n_{\rm sat}$ is the number density of satellites at some altitude, $A_{\rm col}$ is the encounter/collision cross-section of the satellites, $\bar{v}_{r}$ is the typical relative collision speed, and $dV$ is the spherical volume element. The factor of 1/2 is to ensure an encounter between two objects of the same population is not counted twice. This encounter/collision rate can naturally be extended to include combinations of various RSOs (e.g., debris-satellite, debris-debris), as well as non-trivial distributions of collisional area. We also assume that the satellites are randomly distributed within a shell.
To demonstrate the consequences of Figure 1, we can integrate Equation 1 over a limited volume. As an illustrative example, let us integrate over the peak in the density distribution centred on 550 km altitude, assuming a uniform spherical shell. We approximate the shell as being 30 km thick from the peak’s width, and assume a possible collision cross section of $A_{\rm col}\approx 300~\rm m^{2}$, corresponding to a close approach of 10 metres. While such a close approach does not guarantee a collision, depending on the orientation of the satellites involved, it could lead to one, and we use the term “collision” moving forward. Indeed, the average collision distance for randomly oriented thin rods of full length $L$ is approximately $0.3L$. As such, thin satellites but with 30 m solar panel spans can have very large effective cross sections, should there be a loss of control of their orientations. From a safety point of view, using an average physical area for the satellite would underestimate the collision cross section, possibly severely. We discuss this further in Methods.
@@
As also discussed in Methods, we use $\bar{v}_{r}\approx 10~\rm km~s^{-1}$. The density appears to peak at $n_{\rm sat}\approx 3\times 10^{-7}~\rm km^{-3}$, but upon closer inspection we find that the density finely oscillates around $n_{\rm sat}\approx 2\times 10^{-7}~\rm km^{-3}$, so we use this latter value as the average density in the shell.
@@
For the given altitude, the single shell collision rate is thus $\Gamma_{\rm ss}\approx 2.2\times 10^{-6}~\rm s^{-1}$ or about $0.09\rm~d^{-1}$. Defining the single shell expectation time between collisions to be $\tau_{\rm ss}=1/\Gamma_{\rm ss}$, we find $\tau_{\rm ss}\approx 11~\rm d$. This means that the timescale for a 50% chance of one or more potential collisions is about $t\approx 7.6\rm~d$ in that shell alone, assuming a Poisson process, i.e. probability $P=1-\exp(-t/\tau_{\rm ss})$. This result, while approximate, emphasizes that the collision probability for satellites on orbit is substantial without active management.
@@
Although spherical shells around Earth represent extremely large volumes and the instantaneous volume occupied by satellites is small, LEO satellites orbit the Earth in approximately 90 minutes depending on the altitude. As a result, these satellites quickly explore their mutual interaction possibilities. Collision avoidance manoeuvres and station-keeping (i.e. manoeuvres to maintain desired altitudes, separations and orbital phasing), are essential in dense satellite shells and successful active management of constellations seems to be the only reason why there has not been a recent major satellite-satellite collision as orbital densities continue to increase. We will revisit the importance of station-keeping further below.
@@
With this in mind, the corresponding encounter time within each spherical shell is $\tau_{1,h}=1/\Gamma_{1,h}$. The blue curve shows this result in Figure 2, using the density distribution from Figure 1. In the densest part of Starlink’s 550 km orbital shell, we expect close approaches ($1~\rm{km}$) every 22 minutes in that shell alone.
@@
The time between <1<1 km close encounters for all of LEO is found by summing over the rates, with $\tau_{1}=\big(\sum_{h}\Gamma_{1,h}\big)^{-1}$. For all objects, Equation 1 gives $\tau_{1}=36$ seconds. This is dominated by encounters involving at least one satellite (i.e. Sat-RSO encounters), with Starlink (the primary operator in orbit) satellites making up the majority of the satellite population. For Sat-RSO encounters, $\tau_{1}\approx 41~\rm{s}$, while for Starlink-RSO, $\tau_{1}\approx 47~\rm{s}$.
According to the most recent SpaceX biannual report, Starlink satellites made 144,404 collision avoidance manoeuvres in the period between 1 Dec 2024 and 31 May 2025 [SpaceX2025], averaging to 41 manoeuvres per satellite per year, or one collision avoidance manoeuvre every 1.8 minutes across the whole megaconstellation. Keeping in mind that our encounter distance does not directly translate to a collision probability threshold for manoeuvres, nor do we take into account the effects of station-keeping, we interpret the Starlink manoeuvre time to indicate that our simulation results and analytic calculations (1.1 min and 47 s, respectively, see Methods and following discussion regarding our simulations) are reasonable for a no-manoeuvre situation.
@@ 2.1 The CRASH Clock
Taking the estimates above one step further, we can approximate the time for close approaches that could give rise to collisions on orbit. We propose this approach as a measure of the stress in orbital space: the CRASH Clock. For this calculation, we must make some assumptions about collision cross sections.
We use our previously calculated RSO number density distributions as described in Methods, but keep the densities separated into categories (i.e. the number density in a given shell is the sum of $n_{\rm debris}$, $n_{\rm sat}$, etc.). Each interaction combination is then assigned a “collision” cross section (described in Methods). For this discussion, we take collisions to become possible whenever the centre-centre distance between two objects is 10 m (satellite-satellite; rocket body-satellite), 5 m (satellite-debris; rocket-body debris), and 10 cm (debris-debris), although different cross sections could be used.
The total collision rate for a given altitude is
$\Gamma_{h}=\bar{v}_{r}V_{h}\sum_{i,j\leq i}\left(1-\frac{1}{2}\delta_{ij}\right)n_{i}n_{j}A_{ij}^{\rm col}$ (2)
@@
for object types $i$ and $j$, shell volume $V_{h}$ and average relative speed $\bar{v}_{r}=\bar{v}_{r}(h)$. Terms with $i=j$ are halved as in Equation 1. The corresponding collision time is found by summing the rates over all of LEO. For our June 2025 catalogue, this is about $\tau_{\rm col}=5.5\rm~d$ – the CRASH Clock value for that date. For comparison, the no-manoeuvre collision time for Starlink-RSO collisions is about $6\rm~d$. Said differently, if all types of manoeuvres were to suddenly stop, and with our assumptions for collision cross-sections, within 24 hours there is a 17% chance of a close approach that could lead to a collision between two catalogued RSOs, with a 15% that the interaction would involve a Starlink satellite, assuming a Poisson process. Such collisions would be catastrophic, causing a major debris-generating event with high likelihood of secondary and tertiary collisions due to high orbital densities and local collision areas. We repeat this calculation using the LEO TLE catalogue from 1 January 2018, providing a reference value from prior to the megaconstellation era. In stark contrast, the CRASH Clock was 164 days, corresponding to less than a 1% chance of a collision happening between two RSOs within 24 hours under no-manoeuvre conditions. Different CRASH Clock values for different cross sections are discussed further in Methods.
@@
The fidelity of our analytic calculations is compared to direct N-body simulations (see Methods). Simulations were initialized using TLEs for two different epochs: 1 January 2018, and 25 June 2025. The red curves in Figure 2 show the average time between conjunctions less than $1\,{\rm km}$ as a function of altitude, which shows very good agreement with the analytical estimate. Recall that the analytic simulations assume random orbital configurations within each shell, while the simulations use the actual orbits. At least for counting close conjunctions in a statistical sense, our assumption of a random distribution does not appear to be a limitation of the analytic approach. The left-hand panel of Figure 2 shows the 2018 distribution, the right-hand shows 2025. We see that the minimum time between conjunctions has decreased by about two orders of magnitude between 2018 and 2025.
@@
Both the simulations and the analytic calculations are intended to represent encounter statistics for a snapshot in time, akin to a steady state system. That is, we make our calculations using only a chosen TLE catalogue epoch and do not incorporate evolution of the objects on orbit through launches, active debris removal, or PMD. We note that our simulations also do not incorporate atmospheric drag, which would serve to clear out debris at lower altitudes, decrease satellite altitudes, and potentially increase the rate at which satellite orbits would randomize without station-keeping efforts in place.
@@
Additional simulation runs were carried out on these populations using a much smaller time-step (see Methods) and a conjunction distance threshold of $100\,{\rm m}$ to search for conjunction events that could give rise to a collision. For the 25 June 2025 TLEs, we see a $100\,{\rm m}$ conjunction frequency of approximately 1 encounter every 42 minutes, within a factor of two of the analytic estimate (expecting approximately 1 encounter every 60 minutes for distances less than $100\,{\rm m}$). For this particular simulation initialization, a conjunction less than 30 metres occurs in the first 3 hours between debris and a satellite. While this is not inherently a collision, it is an extreme encounter. We note that there is some difficult-to-characterize uncertainty in these numbers because we are using only a single simulation initialization. We have summarized all of our results for both TLE catalogue epochs in Table 1.
@@ 3 Discussion
In an effort to better characterize this stress, for both short and long-term consequences, we recommend the CRASH Clock, which is the expectation time for a collision if all satellite manoeuvres were to suddenly stop. It is a KEI calculated using the methods above, and provides an immediate assessment, as well as historical record, of the stress on the orbital environment. It further provides a measure of the typical time between system-wide loss of control in LEO and a possible catastrophic collision.
We find that our simulations and analytic calculation yield consistent results to within a factor of two. For the purpose of determining the CRASH Clock value, we recommend using the analytic method, even though it assumes randomized orbits. There are several reasons for this. First, the analytic approach is an accessible calculation that only requires a TLE catalogue. Second, it avoids differences in the results that could arise from the choice of numerical methods. Third, simulation techniques could suffer from substantial statistical variation, potentially requiring many simulation initializations. Indeed, such variation can already be seen in the simulation results presented here.
Station-keeping and constellation design deserve a special mention. Satellite constellations can be designed to minimize conflicting interactions among their satellites, with station-keeping being done to maintain optimal configurations [see e.g. sk1, sk2, sk3]. We can thus contemplate situations in which we have ideal constellation designs. In such cases, the CRASH Clock value could in practice be modified to take into account the “phase mixing” time of the constellation in question; by phase mixing time, we mean the theoretical time it would take for the constellation’s orbits to randomize in phase. However, we emphasize again that our analytics and simulations are in reasonable agreement, the latter of which takes into account a snapshot of actual constellation configurations. For this reason, we recommend against removing constellations with idealized designs when determining the CRASH Clock value.
Currently, the 10m-5m-10cm CRASH Clock in LEO is 5.5 days. Yet, the probabilistic nature of collisions means they could happen sooner than any given CRASH Clock value (again in the absence of manoeuvres). Recall that our example numerical simulations showed a debris-satellite conjunction of less than 30 metres in the first 3 hours of simulation time.
@@
A CRASH Clock value of less than a week is already a reason for concern, as major solar storms, such as the May 2024 Gannon storm, can have lingering impacts for the satellite population. For example, in the three days of this storm, nearly half of all active LEO satellites manoeuvred due to increased atmospheric drag to maintain their orbits, and the unpredictable drag in addition to bulk manoeuvring made collision assessment during and after the storm very difficult [Parker2024, Berger]. Additionally, in such conditions positional uncertainties can easily become as high as several km [Parker2024_2], making collision avoidance manoeuvres extremely uncertain. Moreover, while the May 2024 storm was the strongest geomagnetic storm in decades, The Great Geomagnetic Storm of 1859 was at least twice as intense [DSTDATA, CLIVER2013]. That event was characterized by two strong storms within a few days of each other, with the latter commonly known as the Carrington Event (September 1859). The storm peaks lasted for several hours, while the storm durations were for a day and several days, respectively [GREEN2006].
The number of collision avoidance manoeuvres made by Starlink has historically been doubling every six months [Pultarova], although this seems to be the result of several factors, including lowering their risk thresholds (i.e., increasing safety practices). Each manoeuvre can create uncertainty in the estimated satellite positions, with one study even finding inaccuracies immediately after the manoeuvre of up to 40 km [Pultarova2]. As the number of required manoeuvres continues to increase, temporary lapses in collision avoidance capabilities, whether that be from inaccurate orbital determination or even a small miscommunication between operators in manoeuvre decision-making, will become increasingly catastrophic in their potential consequences. Indeed, in 2019 an ESA satellite was forced to manoeuvre out of the way of a Starlink satellite when a bug in SpaceX’s alert system prevented them from seeing an increased collision probability [boyle]. Before modern space traffic management policies, insufficient manoeuvre plan information between operators was the main cause that led to the Iridium-Cosmos collision of 2009 [shepperd].
@@
All the above discussion contextualizes how the the CRASH Clock can be used. It is not a strict limit - rather, it measures risk along a spectrum, with short CRASH Clock times representing a dangerous condition, moderate times representing a caution region, and long times indicating a healthy operational orbital environment. There is some degree of subjectivity in the boundaries of these “danger”, “caution”, and “safe” regions, but this makes the CRASH Clock adaptable without changing the underlying definition. As an illustrative example, and recalling our calculations above, our current CRASH Clock value of $\tau_{\rm col}=5.5~\rm d$ corresponds to a 17% probability of one or more potential collisions during a 24 hour period of no manoeuvre conditions. We consider this to be well within the “caution” region. We could further (and arbitrarily) define the “danger” region to be where the CRASH Clock value implies a 50% chance of at least one potential collision within 24 hours. To stay in the “caution” region and below this “danger” threshold, the CRASH Clock value would therefore need to be longer than 1.4 d.
We could adopt smaller cross sections for the CRASH Clock, indicating when collisions are not just possible but probable. The 4.8m-2.4m-10cm Clock (see Methods) is one example, with a value of about 23 d. From a policy standpoint, however, there needs to be discussion concerning whether the Clock threshold should indicate possible or probable collisions. We have highlighted possible collisions not to be alarmists, but to contextualize the stress and demands on space safety.
Collision risk is not the only issue. We are already experiencing disruption of astronomy [Lawrence2022], pollution in the upper atmosphere from increasingly frequent satellite ablation [Murphy2023], and increased ground casualty risks [Wright2025]. By these safety and pollution metrics, it is clear we have already placed substantial stress on LEO, and changes to our approach are required immediately.
We end by acknowledging that many different groups are working and collaborating to address these issues, especially collision risks. Satellite operators in particular are by no means idle in this regard, as it is in their interest to maintain a safe orbital environment and have teams dedicated to space safety and sustainability. The CRASH Clock is intended to be a tool to help with such efforts.